find two consecutive negative integers who product is 90
Answer by mananth(16949)

let the number be x
the other number will be x-1
x(x-1)=90
x^2-x-90=0
x^2-10x+9x-90=0
x(x-10)+9(x-10)=0
(x-10)(x+9)=0
x=-9, the other number is -10